# Model Predictive and Reinforcement Learning Methods for Active Flow Control of an Airfoil with Dual-point Excitation of Plasma Actuators

### v1

@@ 6 Conclusions
In this study, we explored various advanced control strategies for optimizing the aerodynamic performance of an airfoil by managing flow separation using plasma actuators. The concept previously introduced by the authors, dual-point excitation, is utilized to simultaneously excite both the suction side and trailing-edge shear layers using DBD plasma actuators. The applied control methods include adaptive MPC, TDRL, DQL, and integrated signal processing with the DQL method. These techniques were assessed based on their ability to stabilize the lift coefficient at targeted values under different operating conditions. Our findings revealed significant insights into the effectiveness and limitations of these methods in a highly dynamic and non-linear environment.The main findings are summarized as follows:
@@
      The adaptive MPC approach effectively achieved the desired lift coefficient, $C_{l}$, of 1.60 when operating within the system’s physical limitations.
@@
      Instabilities were observed when attempting to push $C_{l}$ towards the upper physical limits, highlighting challenges in controlling the system near the actuator’s maximum capacity.
@@
      Difficulty was noted in distinguishing between the optimal value of two excitation frequencies (100 Hz and 200 Hz) due to near-identical $C_{l}$ values achieved at these frequencies.
@@
      All RL-based methods successfully converged to an optimal $C_{l}$, albeit with different excitation frequencies.

### v2

@@ 6 Conclusions
In this study, we explored various advanced control strategies for optimizing the aerodynamic performance of an airfoil by managing flow separation using plasma actuators. The concept previously introduced by the authors, dual-point excitation, was employed to simultaneously excite both the suction side and trailing-edge shear layers using DBD plasma actuators. The applied control methods include adaptive MPC, TDRL, DQL, and integrated signal processing with the DQL method. These techniques were assessed based on their ability to stabilize the lift coefficient at targeted values under different operating conditions. Our findings revealed insights into the effectiveness and limitations of these methods in a highly dynamic and non-linear environment. The main findings are summarized as follows:
@@
      The adaptive MPC approach effectively achieved the desired lift coefficient, $C_{l}$, of 1.60, which is within the system’s physical limitations.
@@
      Instabilities were observed when attempting to push $C_{l}$ more than the upper physical limits, highlighting challenges in controlling the system near the actuator’s maximum capacity.
@@
      Challenges were noted in distinguishing between the optimal value of two excitation frequencies (100 Hz and 200 Hz) due to near-identical mean $C_{l}$ values achieved at these frequencies.
@@
      All RL-based methods successfully converged to an optimal $C_{l}$, albeit with different excitation frequencies (Fig. 16).
@@
Further, this study offers new insights into active flow control through the implementation of dual-point DBD plasma actuation. The interaction between this actuation strategy and advanced control techniques provides an effective contribution to mitigating flow separation near stall conditions. The improved aerodynamic performance observed in the controlled cases is primarily attributed to the amplification of shear layer instabilities, which enhance flow entrainment and effectively delay separation. Consequently, a larger suction peak and a more favorable pressure gradient are established over the upper surface of the airfoil, leading to a higher lift coefficient.
From a physical perspective, the differences in control performance between MPC and RL can be linked to how each method handles the unsteady flow dynamics near stall, where nonlinear flow features dominate. The adaptive MPC controller maintains good performance up to a threshold ($C_{L}=1.6$); however, beyond this point, it struggles to ensure stability as the plasma actuator approaches its physical limitations. In contrast, the RL controller adapts more effectively to variations in the separated shear layers, successfully suppressing flow separation and enabling sustained, stable lift enhancement under near-stall conditions.